# Explainability of Large Language Models using SMILE: Statistical Model-agnostic Interpretability with Local Explanations

### v1

@@ 3 Problem Definition
This section introduces the fundamental concepts and mathematical framework necessary to understand our approach. We begin by defining key notation and problem settings, followed by a discussion of measuring the impact of perturbations, the surrogate model used for approximation, and a theoretical justification for its fidelity [13, 100, 101]. Finally, we outline a structured workflow for implementing a Wasserstein-based LIME surrogate, ensuring interpretability in the context of large language models [102].
@@ 3.1 Notation and Problem Setting
In this study, we aim to develop a local surrogate model that explains the behaviour of a black-box model $\pi^{(n)}$ when provided with input text prompts. The basic notations are as follows:
  • Input: $x$, representing the original prompt.
  • Perturbations: $\{\hat{x}_{j}\}_{j=1}^{J}$, small modifications of $x$ used to explore the neighbourhood around the original input.
  • Model outputs: $\pi^{(n)}(y\mid x)$, the model’s probability distribution over possible outputs $y$ for the given input $x$.
@@ 3.2 Input-Level Distance
To measure how different each perturbed input $\hat{x}_{j}$ is from the original prompt $x$, we compute the semantic distance using Word Mover’s Distance (WMD) [103]:
$\delta_{x_{j}}:=\mathrm{WMD}(x,\hat{x}_{j}).$ (1)
Equation 1 quantifies the semantic dissimilarity between $x$ and $\hat{x}_{j}$, ensuring that closer perturbations are treated as more relevant.
@@ 3.3 Output-Level Distribution Shift
To assess how the model’s outputs change when the input is perturbed, we compute the Wasserstein distance between the output distributions for $x$ and $\hat{x}_{j}$ [101]:
$\Delta(x,\hat{x}_{j}):=W\left(\pi^{(n)}(\cdot\mid x),\pi^{(n)}(\cdot\mid\hat{x}_{j})\right).$ (2)
Equation 2 captures the magnitude of the change in the model’s behaviour due to input perturbations.
@@ 3.4 Weighting via Input Similarity
Since not all perturbations are equally meaningful, we prioritise those that are semantically closer to the original input. We define a Gaussian-based weighting scheme [13]:
$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right).$ (3)
As shown in Equation 3, perturbations with smaller semantic distances $\delta_{x_{j}}$ receive higher weights, reflecting their greater relevance.
@@ 3.5 Fitting a Local Surrogate
Each perturbed input $\hat{x}_{j}$ is mapped to a feature vector $z_{j}\in\mathbb{R}^{d}$. The surrogate model $h_{\theta}$ adopts a linear form:
$h_{\theta}(z_{j})=\theta_{0}+\theta^{\top}z_{j}.$ (4)
The parameters $\theta$ are learned by minimising the weighted squared error between the surrogate’s predictions and the observed distribution shifts:
$\min_{\theta}\sum_{j}w_{j}\left(h_{\theta}(z_{j})-\Delta(x,\hat{x}_{j})\right)^{2}.$ (5)
Equation 5 ensures that the surrogate closely approximates the black-box model’s behaviour within the local neighbourhood around $x$.
@@ 3.6 Theoretical Justification (Lipschitz Smoothness)
We formally define the shift function:
$f(x,\hat{x}):=W\left(\pi^{(n)}(\cdot\mid x),\pi^{(n)}(\cdot\mid\hat{x})\right).$ (6)
Assuming that $\pi^{(n)}$ is Lipschitz continuous with respect to its input, the shift function $f$ is also Lipschitz [104]:
$|f(x,\hat{x})-f(x^{\prime},\hat{x}^{\prime})|\leq L\cdot\left(\|x-x^{\prime}\|+\|\hat{x}-\hat{x}^{\prime}\|\right),$ (7)
where $L>0$ is a Lipschitz constant (Equation 7). This smoothness justifies the use of a local linear surrogate model to approximate $f$.
@@
The Kantorovich-Rubinstein duality [105] relates the Wasserstein distance and Total Variation (TV) distance:
@@
where:
@@
Equations 8 and 9 indicate that small changes in the input resulting in small TV distances also imply bounded Wasserstein shifts.
@@ 3.7 Linear Surrogates Approximate Lipschitz Functions
Due to the Lipschitz property, any smooth function $f$ can be locally approximated by a linear model in a small neighbourhood $U$:
$f(x^{\prime})\approx f(x)+\nabla f(x)\cdot(x^{\prime}-x).$ (10)
This principle, expressed in Equation 10, underlies the surrogate modelling approach adopted by LIME and related methods [13].
@@ 3.8 Putting It All Together
The complete Wasserstein-based LIME surrogate process consists of the following steps:
@@
    Generate perturbations $\{\hat{x}_{j}\}$ near $x$.
@@
    Compute semantic distances $\delta_{x_{j}}$ (Equation 1) and output-level shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
@@
    Assign weights $w_{j}$ using Equation 3.
@@
    Extract feature vectors $z_{j}$.
@@
    Fit a local surrogate model $h_{\theta}$ by minimising the loss (Equation 5).
@@
    Use $h_{\theta}$ to interpret how input perturbations affect the model’s output distribution.
This framework provides a rigorous, interpretable method for understanding how small changes to an input affect the outputs of complex models, particularly in applications involving large language models [102, 13, 101, 105].
@@ 4 PROPOSED METHOD
The proposed approach enhances the interpretability of large language models for text generation by analysing how specific textual inputs influence the generated outputs. Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behaviour [8, 102].
In the context of classification tasks, Fig. 4 illustrates SMILE, a tool designed to explain model predictions by isolating critical features that drive decision-making [12]. SMILE segments an input (e.g., an image) into meaningful components: “super-pixels” for images or logical sections for text, and creates perturbations by selectively modifying these components. These perturbations reveal how different input elements contribute to the final prediction.
@@
SMILE records the model’s predictions for each perturbed input and calculates similarity scores between the original and modified inputs using metrics such as the Wasserstein distance [100, 101]. These scores are transformed into weights using a kernel function, enabling a balanced weighting scheme. The results are used to train a weighted linear regression model that identifies the most influential input components.
Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as demonstrated in Fig. 5. Instead of perturbing the generated image, as in SMILE, we modify the instruction by including or excluding certain words. We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between these outputs and the one produced by the original text prompt [103]. This distance is a similarity measure, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
@@
Using this enhanced three-step process, which builds on SMILE’s methodology [12], we aim to help users understand how particular elements of text prompts impact text generation. This technique improves user control and predictability in text generation models by providing insightful information about which elements of textual input, such as stylistic instructions or descriptive terms, most impact the final output.
The first step employs a text generation model based on input text. The model receives the original text input and generates a corresponding output; only the input text changes each time the model renders the output. The text generation model is expected to produce responses similar to the original response that vary only based on critical components of the input texts [107].
Different text permutations are utilised from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into individual words. The instruction-based text generation model then generates text outputs based on both the original input and the perturbed texts [108].
@@ 4.1 Text Generating and perturbed prompts
To begin, we generate variations of the original text prompt, known as “perturbed prompts”, to examine how these subtle changes affect the output of a text generation model [112, 113]. The original text is first broken down into individual words. Multiple text versions are then created by selectively including or excluding specific words. Each perturbed prompt is then provided to the text generation model, which produces a unique output for each variation. In this process, the output generated from the original prompt is the reference, or baseline output, against which all other outputs are compared.
For the original text prompt $p_{\text{org}}$, the text generation function $\phi_{\text{gen}}$ generates a corresponding output text $t_{\text{org}}^{\prime}$, as shown in Eq. 11:
@@
Similarly, for each perturbed prompt $p_{\text{pert}}^{i}$ within the set $S_{\text{pert}}$, the text generation model produces a corresponding output text $t_{\text{pert}}^{\prime i}$, according to Eq. 12:
$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in S_{\text{pert}}.$ (12)
Here, $p_{\text{org}}$ represents the original prompt, $p_{\text{pert}}^{i}$ describes a perturbation generated from the original text, and $\phi_{\text{gen}}$ is the text generation model function used to produce outputs $t^{\prime}$ that reveal the influence of prompt variations. The set of all perturbed prompts is denoted by $S_{\text{pert}}$ [13, 108].
@@ 4.2 Creating the interpretable space
In order to map the output space of the text generation model to a one-dimensional space, the Wasserstein distance between each text generated from the perturbed prompts and the original prompt (baseline output) is computed as a measure of similarity [100, 101]. In all cases, the input prompt remains consistent. Word Mover’s Distance (WMD) is employed to extract meaningful embeddings from the texts, allowing the focus to remain on the primary semantic components of the generated text [103]. This reduces noise in the distance estimation, enabling the accurately calculated distance to serve as the output for the linear regression model.
The distance between the outputs generated by the perturbation texts and the original prompt is given by Eq. 13:
$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{p}\right)^{\frac{1}{p}},\quad\forall i\in S_{\text{pert}}.$ (13)
In the above relation, the $\delta_{\text{WMD}}$ function represents the embedding extraction model applied to the texts before calculating the distance [103]. In Eq. 13, $n$ represents the number of features, $p$ denotes the norm order, and $W(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ demonstrates the Wasserstein distance between the outputs produced by the original and perturbed prompts.
@@
The Wasserstein-$p$ metric allows us to adjust WD’s sensitivity to different text variations by fine-tuning the norm order $p$. Smaller $p$ values emphasise local differences, while larger $p$ values capture broader distributional shifts [101]. By testing various $p$-values, we identify the norm that best captures meaningful variations caused by prompt changes, enhancing the reliability and interpretability of the results [101].
When analysing Wasserstein distances, measures with high p-values, indicating non-significant differences, are excluded from the procedure [114]. As described in Algorithm 1, the Bootstrap algorithm calculates these p-values for all relevant ECDF-based distance measures, including WD. The algorithm computes the WD and its associated p-value.
For a univariate example, let X and Y be the inputs. The Bootstrap algorithm performs 1e5 iterations, where XY is the concatenated set of X and Y. In each iteration, two random samples are drawn from XY, and their Wasserstein distance (boostWD) is computed. If boostWD exceeds the observed WD, a counter (bigger) is incremented. After completing all iterations, the p-value (pVal) is calculated as the proportion of iterations where boostWD is greater than the observed WD, as shown in Algorithm 1.
@@ 14             $\texttt{bigger}\leftarrow\texttt{bigger}+1$;
Figure 5: SMILE flowchart for Large Language Models
When generating perturbation texts derived from the original text, the keywords of the original text play a crucial role in maintaining certain levels of similarity and fidelity between the perturbation texts and the source content. This is important when the interpretable model is trained on perturbed texts as input. Therefore, a similarity criterion is used to measure the degree of similarity between each perturbation text and the original, and this value is employed as a sample weight in the interpretable model. Eq. 15 calculates this similarity, and a Gaussian kernel is applied to normalise it into the range $[0,1]$ [103].
$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\,d(w_{k},w_{l})$ (14)
subject to $\sum_{l=1}^{m}T_{kl}=p_{k},\quad\forall k\in S_{\text{org}}$
$\sum_{k=1}^{n}T_{kl}=q_{l},\quad\forall l\in S_{\text{pert}}$
@@
$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\exp\left(-\left(\frac{C_{i}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma^{2}}\right)^{2}\right),\quad\forall i\in S_{\text{pert}}$ (15)
The function $\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ quantifies the semantic distance between the perturbation text and the original. The resulting weight $\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})$ is obtained by applying a Gaussian kernel to this distance [106], allowing the surrogate model to assign higher importance to more semantically similar inputs.

### v2

@@ 3 Problem Definition
This section introduces the core concepts and mathematical framework underlying our approach. We begin by defining the key notation and problem setup, followed by a formal description of how we quantify the impact of input perturbations. We then describe the use of a surrogate model to approximate model behavior locally and provide theoretical justification based on the Wasserstein distance [63, 112, 113]. Lastly, we outline a structured methodology for implementing a Wasserstein-based surrogate framework tailored to large language models [114].
@@ 3.1 Notation and Problem Setting
We aim to construct a local surrogate model that explains the behavior of a black-box large language model (LLM), denoted as $\pi^{(n)}$, in response to a natural language input prompt. The goal is to understand how specific parts of the input affect the model’s output distribution. The following notation is used:
  • Input: $x\in\mathcal{X}$, representing the original input prompt.
  • Perturbations: $\{\hat{x}_{j}\}_{j=1}^{J}$, a set of $J$ perturbed versions of $x$, generated via minor token-level edits.
  • Model Outputs: $\pi^{(n)}(y\mid\hat{x}_{j})$, the model’s probability distribution over output space $\mathcal{Y}$, conditioned on the perturbed input $\hat{x}_{j}$. For the unperturbed input $x$, we write $\pi^{(n)}(y\mid x)$.
@@ 3.2 Input-Level Distance
To quantify the semantic difference between the original prompt $x$ and each perturbed version $\hat{x}_{j}$, we compute the Word Mover’s Distance (WMD) [115]:
$\delta_{x_{j}}=\mathrm{WMD}(x,\hat{x}_{j}),$ (1)
Where $\delta_{x_{j}}\in\mathbb{R}_{\geq 0}$ denotes the semantic distance between the original and perturbed prompts.
@@ 3.3 Output-Level Distribution Shift
To assess the effect of input perturbation on the model’s behavior, we compute the Wasserstein distance between the output distributions of the original and perturbed inputs:
$\Delta(x,\hat{x}_{j})=W\left(\pi^{(n)}(y\mid x),\pi^{(n)}(y\mid\hat{x}_{j})\right),$ (2)
where $y\in\mathcal{Y}$, and $W(p,q)$ denotes the Wasserstein distance between two probability distributions $p$ and $q$.
@@ 3.4 Weighting via Input Similarity
Perturbations that are semantically closer to the original input are weighted more heavily. We use a Gaussian-based weighting scheme [63]:
$w_{j}=\exp\left(-\left(\frac{\delta_{x_{j}}}{\sigma^{2}}\right)^{2}\right),$ (3)
Where $w_{j}\in(0,1]$ denotes the relevance weight of $\hat{x}_{j}$, and $\sigma>0$ is a kernel width parameter that controls the rate of exponential decay.
@@ 3.5 Fitting a Local Surrogate
Each perturbed input $\hat{x}_{j}$ is mapped to a feature vector $z_{j}\in\mathbb{R}^{d}$. We fit a linear surrogate model of the form:
$h_{\theta}(z_{j})=\theta_{0}+\theta^{\top}z_{j},$ (4)
where $\theta_{0}\in\mathbb{R}$ is a bias term and $\theta\in\mathbb{R}^{d}$ is the weight vector. Parameters $\theta$ are estimated by minimizing the weighted squared error.
$\min_{\theta}\sum_{j=1}^{J}w_{j}\left(h_{\theta}(z_{j})-\Delta(x,\hat{x}_{j})\right)^{2},$ (5)
Where the loss ensures local fidelity between the surrogate and the observed output shifts.
@@ 3.6 Theoretical Justification (Lipschitz Smoothness)
We define a shift function that quantifies the model’s response to perturbation:
$f(x_{a},\hat{x}_{a})=W\left(\pi^{(n)}(y\mid x_{a}),\pi^{(n)}(y\mid\hat{x}_{a})\right),$ (6)
If $\pi^{(n)}$ is Lipschitz continuous concerning its input, then the function $f$ is also Lipschitz:
$\left|f(x_{a},\hat{x}_{a})-f(x_{b},\hat{x}_{b})\right|\leq L\times\left(\|x_{a}-x_{b}\|+\|\hat{x}_{a}-\hat{x}_{b}\|\right),\forall a,b\in\{1,\dots,N_{pert}\}$ (7)
The notation $x_{a}$ and $\hat{x}_{a}$ represents a specific input and its corresponding perturbation. The subscript allows us to extend the formulation to multiple input pairs $(x_{i},\hat{x}_{i})$ when analyzing the Lipschitz continuity or stability across perturbations. $N_{pert}$ denotes the total number of perturbations.
[…]
@@
For two distributions $p$ and $q$ over output space $\mathcal{Y}$, the Kantorovich–Rubinstein duality yields:
@@
Where the total variation distance ($TV$) is given by:
@@
This bound highlights that small changes in distribution (in the TV sense) imply bounded Wasserstein distance.
@@ 3.7 Linear Surrogates Approximate Lipschitz Functions
Due to the differentiability of Lipschitz functions, the shift function $f$ can be locally approximated by a first-order Taylor expansion:
$f(x^{\prime})\approx f(x)+\nabla f(x)^{\top}(x^{\prime}-x),$ (10)
where $\nabla f(x)^{\top}(x^{\prime}-x)$ denotes the dot product between the gradient of $f$ and the local displacement in input space.
@@ 3.8 Putting It All Together
The complete Wasserstein-based LIME surrogate procedure consists of:
@@
    Generating perturbations $\{\hat{x}_{j}\}_{j=1}^{J}$ near input prompt $x$.
@@
    Computing semantic distances $\delta_{x_{j}}$ (Equation 1) and output shifts $\Delta(x,\hat{x}_{j})$ (Equation 2).
@@
    Assigning similarity weights $w_{j}$ via Equation 3.
@@
    Extracting feature representations $z_{j}$ for each perturbation.
@@
    Fitting a local surrogate model $h_{\theta}$ by minimizing the loss (Equation 5).
@@
    Using $h_{\theta}$ to interpret the influence of each input feature on model behavior.
This framework offers a principled and interpretable approach to analyzing large language models through localized perturbation and response analysis.
@@ 4 PROPOSED METHOD
The proposed approach, named gSMILE (Generative SMILE), enhances the interpretability of large language models (LLMs) for text generation by analyzing how specific textual inputs influence the generated outputs. Understanding the interaction between input prompts and generated text improves transparency and predictability in the model’s behavior [9, 114].
In the context of classification tasks, Fig. 4 illustrates SMILE, a tool designed to explain model predictions by isolating critical features that drive decision-making [11]. SMILE segments an input (e.g., an image) into meaningful components: “super-pixels” for images or “logical sections” for text, and creates perturbations by selectively modifying these components. These perturbations reveal how different input elements contribute to the final prediction.
@@
SMILE records the model’s predictions for each perturbed input. It calculates distance scores between the original and modified outputs using metrics such as the Wasserstein distance [112, 113]. These distances are transformed into similarity-based weights using a kernel function, enabling a balanced weighting scheme. The results are used to train a weighted linear regression model that identifies the most influential input components.
Inspired by SMILE, we propose a similar method for interpreting text prompt-based text generation models, as demonstrated in Fig. 5. Instead of perturbing the generated image, as in SMILE, we perturb the input prompt by including or excluding certain words. We generate a corresponding text output for each altered text prompt and compute the Wasserstein distance between each generated output and the original output [115]. This distance quantifies the difference in the production, allowing us to identify which words in the text prompt have the most significant impact on the generated text.
@@
Using this enhanced three-step process, which builds on SMILE’s methodology [11], we aim to help users understand how particular elements of text prompts impact text generation. This technique enhances user control and predictability in text generation models by providing insight into which aspects of textual input, such as stylistic instructions or descriptive terms, have the most significant impact on the final output.
The first step utilizes a text generation model that is trained on input text. The model receives the original text input and generates a corresponding output; only the input text changes each time the model renders the output. The text generation model is expected to produce responses similar to the original response, varying only based on the critical components of the input texts [117].
Different text permutations are utilized from the individual words of the input prompt to create various text perturbation blocks, following the breakdown of the prompt into separate words. The instruction-based text generation model then generates text outputs based on both the original input and the perturbed texts [118].
@@ 4.1 Text Generating and Perturbed Prompts
To begin, we generate variations of the original text prompt, referred to as “perturbed prompts,” to analyze how subtle changes in the input affect the output of a text generation model [122, 123]. The original prompt $p_{\text{org}}$ is first tokenized into individual words. From these tokens, multiple perturbed prompts $\{p_{\text{pert}}^{i}\}_{i=1}^{N}$ are constructed by selectively including or excluding specific words. Each perturbed prompt is then passed to the text generation model, which produces a corresponding output. The output generated from the original prompt, $t_{\text{org}}^{\prime}$, serves as the reference (or baseline) for comparison.
For the original text prompt $p_{\text{org}}$, the text generation model $\phi_{\text{gen}}$ produces the output:
@@
Similarly, for each perturbed prompt $p_{\text{pert}}^{i}\in S_{\text{pert}}$, the model generates:
$t_{\text{pert}}^{\prime i}=\phi_{\text{gen}}(p_{\text{pert}}^{i}),\quad\forall i\in\{1,\dots,N\}$ (12)
Here, $\phi_{\text{gen}}$ is the text generation function, and $t_{\text{pert}}^{\prime i}$ denotes the generated output for the $i$-th perturbation. These outputs are used to measure how individual word-level changes in the prompt influence the model’s behavior [63, 118].
@@ 4.2 Creating the Interpretable Space
To represent the output space of the text generation model in a one-dimensional similarity space, we compute the Wasserstein distance between each generated text from a perturbed prompt and the output generated from the original prompt [112, 113]. Throughout this process, the input structure remains unchanged, and only the perturbed words differ across inputs.
To capture the semantic similarity between generated outputs, we use Word Mover’s Distance (WMD) to embed the texts into a feature space [115]. These embeddings enable the focus to remain on high-level semantic content, thereby mitigating lexical noise and enhancing the reliability of the computed distances. The resulting scalar distances are used as the dependent variable in the subsequent regression model.
The Wasserstein distance between each perturbed output $t_{\text{pert}}^{\prime i}$ and the original output $t_{\text{org}}^{\prime}$ is computed as:
$W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})=\left(\frac{1}{n}\sum_{j=1}^{n}\left\|\delta_{\text{WMD}}(t_{\text{org}}^{\prime j})-\delta_{\text{WMD}}(t_{\text{pert}}^{\prime i,j})\right\|^{P}\right)^{\frac{1}{P}},\quad\forall i\in\{1,\dots,N\}$ (13)
In this equation:
  • $\delta_{\text{WMD}}(\cdot)$ denotes the semantic embedding of a generated text,
  • $n$ is the dimensionality of the embedding space,
  • $P$ specifies the norm order used in the Wasserstein metric,
  • and $W_{i}(t_{\text{org}}^{\prime},t_{\text{pert}}^{\prime i})$ is the Wasserstein distance computed for the $i$-th perturbation.
@@
The Wasserstein-$P$ metric enables adjustment of the sensitivity of the Wasserstein Distance (WD) to different types of text variations by tuning the norm order $P$. Smaller values of $P$ emphasize fine-grained, local differences between distributions, while larger values capture broader, more global shifts [113]. By experimenting with various values of $P$, we can select the one that best captures the semantic differences introduced by prompt perturbations, thereby improving both the reliability and interpretability of the analysis [113].
To assess the statistical significance of the computed Wasserstein distances, we apply hypothesis testing using p-values. Distances associated with high p-values, indicating non-significant variation, are excluded from further consideration [124]. As detailed in Algorithm 1, the Bootstrap algorithm [127, 108] is used to estimate these p-values for each observed WD.
For example, let X and Y denote two sets of text-based embeddings. The Bootstrap algorithm performs 1e5 iterations, where XY is the concatenated set of X and Y. In each iteration, two random samples are drawn from XY, and their Wasserstein distance (boostWD) is computed. If boostWD exceeds the original observed WD, a counter (bigger) is incremented. After all iterations, the p-value (pVal) is estimated as the proportion of iterations where boostWD exceeds the observed WD, as outlined in Algorithm 1.
@@ 14             $\texttt{bigger}\leftarrow\texttt{bigger}+1$;
Figure 5: gSMILE flowchart for Large Language Models
When generating perturbation texts from the original prompt, the key terms in the prompt are essential for maintaining semantic similarity between perturbed prompts and the source. To account for this, we define a similarity metric between each perturbed prompt and the original, which is then used as a sample weight in the interpretable model. This similarity is quantified using Word Mover’s Distance (WMD), which calculates the minimum cost to transport the embedded words of one document to match another [115].
The WMD between the original prompt $p_{\text{org}}$ and a perturbed prompt $p_{\text{pert}}^{i}$ is defined as:
$\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})$ $=\min_{T\geq 0}\sum_{k=1}^{n}\sum_{l=1}^{m}T_{kl}\cdot d(w_{k},w_{l})$ (14)
subject to $\sum_{l=1}^{m}T_{kl}=d_{k},\quad\forall k=1,\dots,n$
$\sum_{k=1}^{n}T_{kl}=d^{\prime}_{l},\quad\forall l=1,\dots,m$
@@
Here:
  • $w_{k}$ and $w_{l}$ denote the $k$-th and $l$-th words from the original and perturbed prompts respectively,
  • $d(w_{k},w_{l})$ is the distance (typically cosine or Euclidean) between the embeddings of $w_{k}$ and $w_{l}$,
  • $T_{kl}$ is the flow variable indicating how much of $w_{k}$’s weight is transported to $w_{l}$,
  • $d_{k}$ and $d^{\prime}_{l}$ represent the normalized bag-of-words weights for the original and perturbed texts.
To use this distance as a weighting factor in the regression model, we normalize it using a Gaussian kernel:
$\pi_{i}(p_{\text{org}},p_{\text{pert}}^{i})=\exp\left(-\left(\frac{\text{WMD}(p_{\text{org}},p_{\text{pert}}^{i})}{\sigma}\right)^{2}\right),\quad\forall i\in\{1,\dots,N\}$ (15)
Here, $\sigma$ is a kernel width parameter that controls sensitivity to dissimilar inputs. The resulting similarity score $\pi_{i}$ falls in the range $[0,1]$, assigning higher weights to perturbed prompts that are semantically closer to the original.